# Mirror Descent Algorithms with Nearly Dimension-Independent Rates for Differentially-Private Stochastic Saddle-Point Problems

### v1

@@ 5 Differentially Private Stochastic Convex Optimization
[…]
Lemma 5.1 (Anytime online-to-batch Cutkosky (2019)).
Let $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ be a convex set. For any sequence $\beta_{t}\in\mathbb{R}^{+},g^{t}\in\mathbb{R}^{d_{x}}$, suppose an online learner predicts $x^{t}$ and receives $t$-th loss $\ell_{t}(x)=\langle g^{t},x\rangle$. Define $w^{t}=\sum_{i\in[t]}\frac{\beta_{i}x^{i}}{\sum_{j\in[t]}\beta_{j}}$. Then, for any convex, differentiable $F$ over $\mathcal{X}$ and any $x\in\mathcal{X},$
@@
An advantage of this conversion is that the weighted averages change slowly from one iteration to the next, which is very useful in the smooth case (Zhang et al., 2022b). Using Lemma 5.1 with weights $\beta_{j}=1$ to optimize over $\Delta_{x}$, the average iterates satisfy $\|w^{t}-w^{t-1}\|_{1}=\|x^{t}-w^{t-1}\|_{1}/t\leq 2/t$. Consequently, $P_{w^{t-1}}$ and $P_{w^{t}}$ are (eventually) similar distributions. We exploit this idea in Algorithm 4 by maintaining estimates of the private iterates for a large number of steps, aggressively reducing the number of samples and improving the privacy. Theorem 5.1 provides a faster rate compared to the one we would get with the plain online-to-batch conversion. Proofs in Appendix F.
@@
Let $f(\cdot;z)$ be convex and first-order-smooth. Then, there exists a parameter setting for $T,K,q,\tau$ such that Algorithm 4 run with these as inputs returns $\hat{w}^{T}$ such that $\hat{w}^{T}$ is $(\varepsilon,\delta)$-DP and

### v2

@@ 4 Differentially Private Stochastic Convex Optimization
We show how our techniques can be used for DP-SCO in the $\ell_{1}$ setting. The setup is similar to DP-SSP in the $\ell_{1}$ setting: consider an unkown distribution $\mathcal{D}$ over a sample space $\mathcal{Z}$. We have access to $\mathcal{D}$ only through a sample $S=\{z^{1},...,z^{n}\}\overset{\text{iid}}{\sim}\mathcal{D}$. Consider a loss function $f:\Delta_{x}\times\mathcal{Z}\to\mathbb{R}$, where $\Delta_{x}$ is the standard simplex in $\mathbb{R}^{d_{x}}$. We assume that for all $z\in\mathcal{Z}$, the function $f(\cdot,z)$ is $L_{1}$-smooth and $L_{0}$-Lipschitz w.r.t $\|\cdot\|_{1}$ over $\Delta_{x}$. Let $F_{\mathcal{D}}(x)=\mathbb{E}_{z\sim\mathcal{D}}[f(x;z)]$ be the population loss. We are interested in solving $\min_{x\in\Delta_{x}}F_{\mathcal{D}}(x)$ with $(\varepsilon,\delta)$-DP algorithms. Many ideas used in the previous section for the SSP problem will be useful for here as well. We observe that guarantees for (entropic) stochastic mirror descent can be obtained by applying online-to-batch conversions on exponentiated gradient descent [orabona_online_learning]. By contrast, we will next apply anytime online-to-batch conversions to exponentiated gradient descent [anytime-online-to-batch]. The version we provide below for this result is slightly different than [anytime-online-to-batch] (see [online_to_batch] for a proof).
Lemma 4.1 (Anytime online-to-batch [anytime-online-to-batch, online_to_batch]).
Let $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ be a convex set. For any sequence $\beta_{t}\in\mathbb{R}^{+},g^{t}\in\mathbb{R}^{d_{x}}$, suppose an online learner predicts $x^{t}$ and receives $t$-th loss $\ell_{t}(x)=\langle g^{t},x\rangle$. Define $w^{t}=\sum_{i\in[t]}\frac{\beta_{i}x^{i}}{\sum_{j\in[t]}\beta_{j}}$ and $\operatorname{Regret}_{T}(x)=\sum_{t\in[T]}\langle g_{t},x_{t}-x\rangle$. Then, for any convex, differentiable $F$ over $\mathcal{X}$ and any $x\in\mathcal{X},$
@@
An advantage of this conversion is that the weighted averages $w^{t}$ change slowly from one iteration to the next, which is very useful in the smooth case [online_to_batch]. Using Lemma 4.1 with weights $\beta_{j}=1$ to optimize over $\Delta_{x}$, the average iterates satisfy $\|w^{t}-w^{t-1}\|_{1}=\|x^{t}-w^{t-1}\|_{1}/t\leq 2/t$. Consequently, $P_{w^{t-1}}$ and $P_{w^{t}}$ are progressively closer distributions. We exploit this idea in Algorithm 5 by maintaining estimates of the private iterates for a large number of steps, aggressively reducing the number of sampled vertices and improving the privacy. Theorem 4.2 provides a faster rate compared to the one we would get with the plain online-to-batch conversion.
@@
Let $0<\delta<1$, $0<\varepsilon<8\log(1/\delta)$. Let $f(\cdot;z)$ be convex and first-order-smooth. Then, there exists a parameter setting for $T,K,q,\tau$ such that Algorithm 5 run with these as inputs returns $\hat{w}^{T}$ such that $\hat{w}^{T}$ is $(\varepsilon,\delta)$-DP and